# General effective reduction theory of integral polynomials of given non-zero discriminant and its applications

### v1

@@ Abstract.
We give a survey on the general effective reduction theory of integral polynomials and its applications. We concentrate on results providing the finiteness for the number of ‘$\mathbb{Z}$-equivalence classes’ and ‘$GL_{2}(\mathbb{Z})$-equivalence classes’ of polynomials of given discriminant. We present the effective finiteness results of Lagrange from 1773 and Hermite from 1848, 1851 for quadratic resp. cubic polynomials. Then we formulate the general ineffective finiteness result of Birch and Merriman from 1972, the general effective finiteness theorems of Győry from 1973, obtained independently, and of Evertse and Győry from 1991, and a result of Hermite from 1857 not discussed in the literature before 2023 for a reason explained below. We briefly outline our effective proofs which depend on Győry’s effective results on unit equations, whose proofs involve Baker’s theory of logarithmic forms. Then we focus on our recent joint paper with Bhargava, Remete and Swaminathan from 2023, where Hermite’s finiteness result from 1857 involving ‘Hermite equivalence classes’ is compared with the above-mentioned modern results, using the two classical equivalences, and where it is confirmed that Hermite’s result from 1857 is much weaker than the modern results mentioned. Since 1973, the results of Győry from 1973 and Evertse and Győry from 1991 together established a general effective reduction theory of integral polynomials with given non-zero discriminant which resulted in many significant consequences and applications, including Győry’s effective finiteness theorems from the 1970’s on monogenic orders and number fields. For generalizations and further applications we refer to our monograph [J.-H. Evertse and K. Győry, Discriminant Equations in Diophantine Number Theory, Camb. New Math. Monogr 32, Cambridge University Press, 2017] and Sections 58 of the present paper. In the Appendix we discuss related topics not strictly belonging to the reduction theory of integral polynomials considered.
@@
Keywords and Phrases: polynomials, discriminants, $\mathbb{Z}$-equivalence, $GL_{2}(\mathbb{Z})$-equivalence, Hermite equivalence, power integral bases, monogenic number fields
@@ 1.1. Preliminaries
Two polynomials $f,g\in\mathbb{Z}[X]$ of degree $n$ are called $\mathbb{Z}$-equivalent if $g(X)=f(X+a)$ or $g(X)=(-1)^{n}f(-X+a)$ for some $a\in\mathbb{Z}$, and $GL_{2}(\mathbb{Z})$-equivalent if $g(X)=\pm(cX+d)^{n}f\big{(}\frac{aX+b}{cX+d}\big{)}$ for some matrix $\big{(}\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big{)}\in GL_{2}(\mathbb{Z})$, i.e., $a,b,c,d\in\mathbb{Z}$ and $ad-bc=\pm 1$. Clearly, $\mathbb{Z}$-equivalence implies $GL_{2}(\mathbb{Z})$-equivalence. Polynomials that are $\mathbb{Z}$-equivalent to a monic polynomial are also monic.
@@
$f=a_{0}X^{n}+\cdots+a_{n}=a_{0}\prod_{i=1}^{n}(X-\alpha_{i})$
@@
This is a homogeneous polynomial of degree $22n-$ 2 in $\mathbb{Z}[a_{0},\ldots,a_{n}]$; thus, if $f\in\mathbb{Z}[X]$ then $D(f)\in\mathbb{Z}$. As one may easily verify, polynomials that are $\mathbb{Z}$-equivalent or $GL_{2}(\mathbb{Z})$-equivalent have the same discriminant.
@@
Roughly speaking, reduction theory of polynomials entails to find, for a given polynomial $f\in\mathbb{Z}[X]$, another polynomial $g\in\mathbb{Z}[X]$ that is $GL_{2}(\mathbb{Z})$-equivalent (or $\mathbb{Z}$-equivalent in the monic case) to $f$ and whose coefficients have as small as possible absolute values. In this paper, we focus on results in which the height $H(g)$ of $g$ is bounded above in terms of $|D(f)|$, i.e., on reduction theory for polynomials of given discriminant. Such results imply that up to $GL_{2}(\mathbb{Z})$-equivalence (resp. $\mathbb{Z}$-equivalence if we restrict ourselves to monic polynomials) there are only finitely many polynomials $f\in\mathbb{Z}[X]$ of degree $n$ and given discriminant $D\not=0$.
The results on reduction theory for univariate polynomials can be translated immediately into similar results for binary forms. We decided to formulate our results in terms of univariate polynomials for convenience of presentation.
1.2. Brief summary of results
@@
The reduction theory of integral polynomials of given non-zero discriminant was initiated by Lagrange (1773). For quadratic polynomials he proved that up to the classical $GL_{2}(\mathbb{Z})$-equivalence, resp. $\mathbb{Z}$-equivalence (monic case) there are only finitely many quadratic polynomials in $\mathbb{Z}[X]$ of given discriminant. Lagrange’s result is effective in the sense that one can effectively determine the polynomials.
@@
Hermite (1848, 1851) introduced a reduction theory for polynomials of arbitrary degree but using another invariant instead of the discriminant, which for cubic polynomials gives the analogue of Lagrange’s result. Hermite was apparently interested to extend this to the general case, i.e., that for every $n\geq 4$ and $D\not=0$ there are up to $GL_{2}(\mathbb{Z})$-equivalence only finitely many polynomials $f\in\mathbb{Z}[X]$ of degree $n$ and discriminant $D$. In Hermite (1857) he introduced a new equivalence relation (called by us ‘Hermite equivalence’, see Section 3) and proved in an ineffective way a finiteness result on the corresponding equivalence classes of integral polynomials of degree $n$ and discriminant $D$. But he did not compare his equivalence relation to the classical equivalence relations, i.e., to $GL_{2}(\mathbb{Z})$-equivalence and $\mathbb{Z}$-equivalence. The result of Hermite (1857) does not appear to have been studied in the literature until the excellent book of Narkiewicz (2018), where Hermite equivalence was mixed up with the classical equivalence relations.
@@
Hermite’s apparent goal, i.e., the finiteness result with $GL_{2}(\mathbb{Z})$-equivalence instead of Hermite equivalence, was finally achieved more than a century later by Birch and Merriman (1972) for arbitrary polynomials in an ineffective form and independently, for monic polynomials and in a more precise and effective form by Győry (1973). The result of Birch and Merriman was subsequently made effective by Evertse and Győry (1991). More precisely, Győry (1973) and Evertse and Győry (1991) proved that there exists an effectively computable number $c(D)$ depending only on $D$ such that every $f\in\mathbb{Z}[X]$ of discriminant $D\not=0$ is $GL_{2}(\mathbb{Z})$-equivalent (and even $\mathbb{Z}$-equivalent in the monic case) to a polynomial $g$ whose coefficients have absolute values bounded above by $c(D)$. These results heavily depend on effective finiteness results for unit equations $ax+by=1$ with solutions $x,y$ from the unit group of the ring of integers of a number field, which were derived in turn using Baker’s theory of logarithmic forms. This solved the old problem of Hermite (1857) mentioned above in an effective way, and further resulted in many significant consequences and applications. For example, in the 1970’s, Győry deduced from his paper from 1973 the first general effective algorithm that decides monogenicity and existence of power integral bases of number fields, and in fact finds all power integral bases. For later applications and generalizations we refer to the monograph Evertse and Győry (2017) and Sections 58 of the present paper.
@@
In our recent paper BEGyRS (2023) with Bhargava, Remete and Swaminathan we provided a thorough treatment of the notion of Hermite equivalence, and proved that $\mathbb{Z}$-equivalence and $GL_{2}(\mathbb{Z})$-equivalence are much more precise than Hermite equivalence. This confirmed that Hermite’s result from 1857 was much weaker than those of Birch and Merriman, Győry, and that of Evertse and Győry mentioned above. It should of course be mentioned that unlike the last authors, Hermite didn’t have the powerful Baker’s theory of logarithmic forms at his disposal.
In the remainder of the paper we discuss in more detail the results of Birch and Merriman (1972), Győry (1973), Evertse and Győry (1991), and those from the paper BEGyRS (2023). In the Appendix we go into some related topics not strictly belonging to reduction theory of integral polynomials. In a later, yet to be written extended version of our present paper we shall publish Section B from the Appendix, which deals with so-called rationally monogenic orders, in a more detailed form.
@@ 2. Reduction theory of integral quadratic and cubic polynomials of given non-zero discriminant
As we mentioned, Lagrange (1773) was the first to develop a reduction theory for binary quadratic forms with integral coefficients. His theory was made more precise by Gauss (1801). For integral polynomials, their theories imply the following. Recall that the height $H(g)$ of a polynomial with integer coefficients is the maximum of the absolute values of its coefficients.
@@
The proofs of Delone and Nagell of Theorem 2.5 were both ineffective, in that they did not provide a method to determine the polynomials. In fact, these proofs were based on a classical ineffective finiteness theorem of Thue (1909) on Thue equations, i.e. on equations of the form $F(x,y)=m$, $x,y\in\mathbb{Z}$, where $F\in\mathbb{Z}[X,Y]$ is a binary form of degree $\geq 3$ with discriminant $\neq 0$ and $m$ is a non-zero integer. In some concrete cases Delone and Faddeev (1940) made effective Theorem 2.5, and posed the problem to make it effective for any irreducible cubic monic polynomial. An effective version of Theorem 2.5 follows from the famous effective result of Baker (1968) on Thue equations.
3. Hermite’s attempt (1857) to extend the previous reduction results to polynomials of arbitrary degree
@@ 3.1. $\mathbf{GL_{n}(\mathbb{Z})}$-equivalence of decomposable forms
Hermite tried to extend his theorem (1851) on cubic integral binary forms resp. polynomials to the case of arbitrary degree $n\geq 4$, but without success. Instead, he proved a finiteness theorem with a weaker equivalence, see Theorem 3.2 below. Hermite’s notion of equivalence (called by us ‘Hermite equivalence’) is based on an equivalence relation for certain decomposable forms.
@@
Let $GL_{n}(\mathbb{Z})$ denote the multiplicative group of $n\times n$ integer matrices of determinant ±$\pm 1$. Two decomposable forms $F,G$ as above are called $GL_{n}(\mathbb{Z})$-equivalent if
$G(\mathbf{X})=\pm F(U\mathbf{X})\text{{ for some }}U\in GL_{n}(\mathbb{Z}),$
@@ 3.2. Hermite equivalence of polynomials and Hermite’s finiteness theorem
$[f](\mathbf{X}):=a_{0}^{n-1}\prod_{i=1}^{n}(X_{1}+\alpha_{i}X_{2}+\cdots+\alpha_{i}^{n-1}X_{n})\in\mathbb{Z}[X_{1},\ldots,X_{n}].$
@@
Since $\mathbb{Z}$-equivalence implies $GL_{2}(\mathbb{Z})$-equivalence, it suffices to prove Proposition 3.3 for $GL_{2}(\mathbb{Z})$-equivalence. We recall the proof from BEGyRS (2023).
@@
$[g](\mathbf{X})$ =±$=\pm a_{0}^{n-1}\prod_{i=1}^{n}\langle\mathbf{b}_{i},\mathbf{X}\rangle\text{, where }\mathbf{b}_{i}=(\beta_{i}^{n-1},\beta_{i}^{n-2}\gamma_{i},\ldots,\gamma_{i}^{n-1})^{T}.$
@@
=±$=\pm c^{n-1}\prod_{i=1}^{n}\langle\mathbf{a}_{i},t(A)^{T}\mathbf{X}\rangle=\pm[f](t(A)^{T}\mathbf{X}),$
@@
In BEGyRS (2023) we gave, for every $n\geq 4$ and both for the nonmonic and for the monic case, infinite collections of polynomials in $\mathbb{Z}[X]$ with degree $n$ that are Hermite equivalent but not $GL_{2}(\mathbb{Z})$-equivalent. More precisely we proved the following.
@@
    There exist infinitely many Hermite equivalence classes of properly nonmonic [^1] That is, not $GL_{2}(\mathbb{Z})$-equivalent to any monic polynomial irreducible polynomials of degree $n$ that split into more than one $GL_{2}(\mathbb{Z})$-equivalence class.
@@
We proved Theorem 3.4 simultaneously for the cases (i) and (ii). We constructed, for every integer $n\geq 4$, an infinite parametric family of pairs $(f_{t,c}^{(n)},g_{t,c}^{(n)})$ of primitive  [^2] An integral polynomial is called primitive if its coefficients have greatest common divisor $1$, irreducible polynomials $f_{t,c}^{(n)}$, $g_{t,c}^{(n)}$ of degree $n$, where $c$ runs through $1$ and an infinite set of primes, and $t$ runs through an infinite set of primes with $t\neq c$ with the following properties:
(3.2) $f_{t,c}^{(n)},\,g_{t,c}^{(n)}$ have leading coefficient $c$ and are properly
nonmonic if $c>1$;
(3.3) $f_{t,c}^{(n)},\,g_{t,c}^{(n)}$ are Hermite equivalent;
(3.4) $f_{t,c}^{(n)},\,g_{t,c}^{(n)}$ are not $GL_{2}(\mathbb{Z})$-equivalent;
(3.5) the pairs $(f_{t,c}^{(n)},g_{t,c}^{(n)})$ lie in different Hermite equivalence
classes.
The main steps of the proof are as follows. From the construction of $f_{t,c}^{(n)}$ and $g_{t,c}^{(n)}$ it is easy to show that (3.2) and (3.3) hold. The proof of (3.4) is more complicated. It requires the use of an irreducibility theorem of Dumas (1906), Chebotarev’s density theorem, and Dirichlet’s theorem on primes in arithmetic progressions. Finally, $f_{t,c}^{(n)}$ is so chosen that if we fix $n,c$ and let $t\to\infty$ then the absolute value of the discriminant of $f_{t,c}^{(n)}$ tends to $\infty$. Since Hermite equivalent polynomials have the same discriminant, the pairs $(f_{t,c}^{(n)},g_{t,c}^{(n)})$ lie in infinitely many different Hermite equivalence classes, that is, (3.5) follows.
@@
We note that in our paper BEGyRS (2023) it turned out that the Hermite equivalence class of a polynomial has a very natural interpretation in terms of the so-called invariant ring and invariant ideal associated with the polynomial, see Appendix B for more details. This fact turned out to be important in the above proofs.
Proposition 3.3 and Theorem 3.4 imply that $GL_{2}(\mathbb{Z})$-equivalence, resp. $\mathbb{Z}$-equivalence are stronger than Hermite equivalence, and hence that Hermite’s Theorem 3.2 is much weaker than the most significant results of this area presented in Section 4 below.
4. Reduction theory of integral polynomials of given non-zero discriminant: the general case
As we mentioned in the Introduction, the breakthroughs in the reduction theory due to Birch and Merriman (1972), Győry (1973), and Evertse and Győry (1991) settled the old problem of Hermite (1857), to prove that for every given $n\geq 2$ and $D\not=0$ there are up to $GL_{2}(\mathbb{Z})$-equivalence only finitely many polynomials $f\in\mathbb{Z}[X]$ of degree $n$ and discriminant $D$, and to determine these effectively. We state the results in more detail.
4.1. The theorems of Birch and Merriman (1972), Győry (1973) and Evertse and Győry (1991)

### v2

# Effective reduction theory of integral polynomials of given discriminant, and related topics

@@ Abstract.
For polynomials in $\mathbb{Z}[X]$, the classical $\mathbb{Z}$-equivalence (monic case) and $GL_{2}(\mathbb{Z})$-equivalence preserve the discriminant as an invariant. The effective reduction theory for polynomials of given degree and discriminant consists of results that give, for a given polynomial $f\in\mathbb{Z}[X]$, a $\mathbb{Z}$-equivalent (in the monic case) or $GL_{2}(\mathbb{Z})$-equivalent polynomial $g$ whose coefficients are effectively bounded above in terms of only the degree and discriminant of $f$. We discuss the classical results of this type of Lagrange (1773) and Hermite (1851) on quadratic and cubic polynomials, the general ineffective theorem of Birch and Merriman (1972), the general effective theorem of Győry (1973) for monic polynomials, obtained independently, and that of Evertse and Győry (1991) for arbitrary polynomials. The proofs of these two effective theorems use Győry’s effective results on unit equations, which were proved by means of Baker’s effective theory of logarithmic forms. Later Evertse, Győry and others obtained several applications and generalizations; see the book Evertse and Győry (2017). In his long-forgotten paper Hermite (1857), Hermite attempted to extend the above results of Lagrange and Hermite to polynomials of arbitrary degree. However, as was pointed out in our joint work BEGyRS (2023) with Bhargava, Remete and Swaminathan, Hermite (1857) proved an important result but with a weaker equivalence only. Thus, it was only by the above mentioned theorems of Győry (1973) and Evertse and Győry (1991) that Hermite’s problem from 1857 was settled in full generality. This and many other recent results inspired us to write this survey paper on the subject. We present here several older and recent generalizations and applications of the effective reduction theory, e.g., to monogenic number fields and monogenic and rationally monogenic orders. We also give an overview of bounds on the number of times a given order is monogenic or rationally monogenic. In the Appendix we discuss further related topics not strictly belonging to the reduction theory of integral polynomials.
@@
Keywords and Phrases: polynomials, discriminants, $\mathbb{Z}$-equivalence, $GL_{2}(\mathbb{Z})$-equivalence, Hermite equivalence, power integral bases, monogenic number fields, monogenic and rationally monogenic orders
@@ 1.1. Preliminaries
Two polynomials $f,g\in\mathbb{Z}[X]$ of degree $n\geq 2$ are called $\mathbb{Z}$-equivalent if
$g(X)=f(X+a)$ or $g(X)=(-1)^{n}f(-X+a)$ for some $a\in\mathbb{Z}$,
and $GL_{2}(\mathbb{Z})$-equivalent if
$g(X)=\pm(cX+d)^{n}f\big(\mbox{\large{$\textstyle{\frac{aX+b}{cX+d}}$}}\big)$ for some matrix $\big(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big)\in GL_{2}(\mathbb{Z})$,
i.e., $a,b,c,d\in\mathbb{Z}$ and $ad-bc=\pm 1$. Clearly, $\mathbb{Z}$-equivalence implies $GL_{2}(\mathbb{Z})$-equivalence. Polynomials that are $\mathbb{Z}$-equivalent to a monic polynomial are also monic.
@@
$f=a_{0}X^{n}+\cdots+a_{n}=a_{0}\prod_{i=1}^{n}(X-\alpha_{i}),\ \ \text{with }a_{0}\not=0$
@@
This is a homogeneous polynomial of degree 2n22n-2 in $\mathbb{Z}[a_{0},\ldots,a_{n}]$; thus, if $f\in\mathbb{Z}[X]$ then $D(f)\in\mathbb{Z}$. As one may easily verify, polynomials that are $\mathbb{Z}$-equivalent or $GL_{2}(\mathbb{Z})$-equivalent have the same discriminant.
@@
An invariant is a function $\mathbb{Z}[X]\to\mathbb{R}$ that assumes the same value at $GL_{2}(\mathbb{Z})$-equivalent polynomials. In general, reduction theory of polynomials is about results of the following type: given a set of invariants, $I_{1},\ldots,I_{t}$, say, there exists for any $f\in\mathbb{Z}[X]$ a polynomial $g\in\mathbb{Z}[X]$ that is $GL_{2}(\mathbb{Z})$-equivalent (or $\mathbb{Z}$-equivalent in the monic case) to $f$ and whose coefficients are bounded in terms of $I_{1}(f),\ldots,I_{t}(f)$. In this paper, we focus on results in which the height $H(g)$ of $g$ is bounded above in terms of $\deg f$ and $|D(f)|$, i.e., on reduction theory for polynomials of given degree and given discriminant. Such results imply that up to $GL_{2}(\mathbb{Z})$-equivalence (resp. $\mathbb{Z}$-equivalence if we restrict ourselves to monic polynomials) there are only finitely many polynomials $f\in\mathbb{Z}[X]$ of degree $n$ and given discriminant $D\not=0$.
In fact, most of the literature deals with reduction theory of binary forms of given discriminant. Recall that any binary form $F(X,Y)\in\mathbb{Z}[X,Y]$ can be factored as $\prod_{i=1}^{n}(\alpha_{i}X-\beta_{i}Y)$ with algebraic $\alpha_{i},\beta_{i}$, and that its discriminant is $D(F):=\prod_{1\leq i<j\leq n}(\alpha_{i}\beta_{j}-\alpha_{j}\beta_{i})^{2}$. Two binary forms $F,G\in\mathbb{Z}[X,Y]$ are called $GL_{2}(\mathbb{Z})$-equivalent if $G(X,Y)=\pm F(aX+bY,cX+dY)$ for some $\big(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\big)\in GL_{2}(\mathbb{Z})$, and clearly, $GL_{2}(\mathbb{Z})$-equivalent binary forms have the same discriminant. The results on reduction theory for binary forms $F$ can be translated immediately into similar results for univariate polynomials $f$ and vice-versa, using the correspondence $f(X)=F(X,1)$, $F(X,Y)=Y^{\deg f}f(X/Y)$. As in our joint paper BEGyRS (2023) with Bhargava, Remete and Swaminathan, we formulate our results in terms of univariate polynomials for convenience of presentation.
For definitions of effectively given concepts, structures and effective determination, computation, one can consult e.g. the corresponding sections of our books Evertse and Győry (2015, 2017, 2022).
1.2. Summary
@@
Lagrange (1773) developed a reduction theory of integral binary quadratic forms of given discriminant, which can be translated immediately into a reduction theory for quadratic polynomials of given non-zero discriminant. His results imply that up to the classical $GL_{2}(\mathbb{Z})$-equivalence, resp. $\mathbb{Z}$-equivalence (monic case) there are only finitely many quadratic polynomials in $\mathbb{Z}[X]$ of given discriminant. Lagrange’s result is effective in the sense that one can effectively determine the reduced polynomials. This was later made more precise by Gauss (1801).
@@
Hermite (1848, 1851) introduced a reduction theory for binary forms, or equivalently univariate polynomials of arbitrary degree but using another invariant instead of the discriminant. In the case of cubic polynomials, Hermite’s invariant is up to a constant a power of the absolute value of the discriminant. Thus, Hermite’s reduction theory implies that up to $GL_{2}(\mathbb{Z})$-equivalence there are only finitely many cubic polynomials in $\mathbb{Z}[X]$ of given discriminant. Hermite was apparently interested to extend this to polynomials of arbitrary degree $n\geq 4$. In Hermite (1857) he introduced a new equivalence relation (called by us ‘Hermite equivalence’, see Section 3) and proved in an ineffective way a finiteness result on the corresponding equivalence classes of integral polynomials of degree $n$ and discriminant $D$. But he did not compare his equivalence relation to the classical equivalence relations, i.e., to $GL_{2}(\mathbb{Z})$-equivalence and $\mathbb{Z}$-equivalence. The result of Hermite (1857) does not appear to have been studied in the literature until the excellent book of Narkiewicz (2018), where Hermite equivalence was confused with the classical equivalence relations.
@@
Hermite’s apparent goal, i.e., the finiteness result with $GL_{2}(\mathbb{Z})$-equivalence instead of Hermite equivalence, was finally achieved more than a century later by Birch and Merriman (1972) for arbitrary polynomials in an ineffective form and independently, for monic polynomials and in a more precise and effective form by Győry (1973). The general result of Birch and Merriman was subsequently made effective by Evertse and Győry (1991a). More precisely, Győry (1973) and Evertse and Győry (1991a) proved that there exists an effectively computable number $c(n,D)$ depending only on $n$ and $D$ such that every $f\in\mathbb{Z}[X]$ of degree $n$ and discriminant $D\not=0$ is $GL_{2}(\mathbb{Z})$-equivalent (and even $\mathbb{Z}$-equivalent in the monic case) to a polynomial $g$ with height
(1.1) $H(g)\leq c(n,D).$
These results heavily depend on effective finiteness results for unit equations $ax+by=1$ with solutions $x,y$ from the unit group of the ring of integers of a number field, which were derived in turn using Baker’s theory of logarithmic forms. This solved the old problem of Hermite (1857) mentioned above in an effective way, and further resulted in many significant consequences and applications. For example, in the 1970’s, Győry deduced from his paper from 1973 the first general effective algorithm that decides monogenicity and existence of power integral bases of number fields, and in fact finds all power integral bases. For later applications and generalizations we refer to the monograph Evertse and Győry (2017) and Sections 49 of the present paper.
@@
In our recent paper BEGyRS (2023) with Bhargava, Remete and Swaminathan we provided a thorough treatment of the notion of Hermite equivalence, and proved that $\mathbb{Z}$-equivalence and $GL_{2}(\mathbb{Z})$-equivalence are much more precise than Hermite equivalence. This confirmed that Hermite’s result from 1857 was weaker than those of Birch and Merriman, Győry, and that of Evertse and Győry mentioned above. It should of course be mentioned that unlike the last authors, Hermite didn’t have the powerful Baker’s theory of logarithmic forms and its application to unit equations at his disposal.
In Section 2 we briefly recall the reduction theory of quadratic and cubic polynomials of given non-zero discriminant. In Section 3, following BEGyRS (2023), we deal with Hermite equivalence and compare it with $\mathbb{Z}$-equivalence and $GL_{2}(\mathbb{Z})$-equivalence. In Section 4 we discuss in more detail the general results of Birch and Merriman (1972), Győry (1973), Evertse and Győry (1991a), and those from the paper BEGyRS (2023). We present the best known effective height estimates for the solutions of unit equations and $S$-unit equations. We sketch how to deduce results of the type (1.1). An important part of Section 4 is Subsection 4.7, which gives much stronger conjectural upper bounds for the height of $g$. These bounds follow from the abc-conjecture and related conjectures. This is partly joint work with Rafael von Känel. In Section 5 we present some consequences in algebraic number theory. In particular, we give an overview of effective finiteness results concerning algebraic numbers of given discriminant, resp. given index, and index form equations. Further, we deduce applications to monogenic number fields and orders, and also generalizations to so-called rationally monogenic orders. In Section 6 we discuss practical algorithms for solving index form equations, i.e., determining all power integral bases in number fields of degree $\leq 6$. In Section 7 we give applications to canonical number systems in number fields and orders, and in Section 8 to some classical Diophantine equations. Section 9 gives a brief overview of generalizations, among others to the number field and $\mathfrak{p}$-adic case, and to results where the ground ring is of characteristic 0 and finitely generated as a $\mathbb{Z}$-algebra. In Section 10 we give an overview of results concerning multiply monogenic and rationally monogenic orders, where we present uniform upper bounds for the multiplicity of (rational) monogenicity of orders, depending only on the degree of the underlying number field. In the Appendix we briefly discuss related topics not strictly belonging to reduction theory of integral polynomials, in particular statistical results for monogenic and rationally monogenic number fields, and Hasse’s problem to give an arithmetic characterization of the monogenic number fields.
Remark. Since the 1970’s, the reduction theory of integral polynomials of given discriminant has been constantly developing, with a growing number of results and applications. Except for Section 2, the other sections contain results from this period. We propose some problems, whose solutions would yield considerable progress in the reduction theory.
Acknowledgments. We are very much indebted to Professor Rafael von Känel for his important contributions to Subsection 4.7, and Professor Attila Pethő for his useful comments on Section 7. We are very grateful to Dr. Csanád Bertók for typing a substantial part of the manuscript. The second named author was supported in part from the Austrian-Hungarian joint project ANN130909 (FWF-NKFIH) and from NKFIH 150284.
@@ 2. Reduction theory of integral quadratic and cubic polynomials of given non-zero discriminant
As we mentioned, Lagrange (1773) was the first to develop a reduction theory for binary quadratic forms with integral coefficients. His theory was made more precise by Gauss (1801). For integral polynomials, their theories imply the following. Recall that the height $H(g)$ of a polynomial with integral coefficients is the maximum of the absolute values of its coefficients.
@@
The proofs of Delone and Nagell of Theorem 2.5 were both ineffective, in that they did not provide a method to determine the polynomials. In fact, these proofs were based on a classical ineffective finiteness theorem of Thue (1909) on Thue equations, i.e. on equations of the form $F(x,y)=m$, $x,y\in\mathbb{Z}$, where $F\in\mathbb{Z}[X,Y]$ is an irreducible binary form of degree $\geq 3$ and $m$ is an integer. In some concrete cases Delone and Faddeev (1940) made effective Theorem 2.5, and posed the problem to make it effective for any irreducible cubic monic polynomial. An effective version of Theorem 2.5 follows from the famous effective result of Baker (1968b) on Thue equations.
3. Hermite’s attempt (1857) to extend the reduction results of polynomials of degree $\leq 3$ to polynomials of arbitrary degree
@@ 3.1. $\mathbf{GL_{n}(\mathbb{Z})}$-equivalence of decomposable forms
Hermite tried to extend his theorem (1851) on cubic integral binary forms resp. polynomials to the case of any degree $n\geq 4$, but without success. Instead, he proved a finiteness theorem with a weaker equivalence, see Theorem 3.2 below. Hermite’s notion of equivalence (called by us ‘Hermite equivalence’) is based on an equivalence relation for certain decomposable forms.
@@
Let $GL_{n}(\mathbb{Z})$ denote the multiplicative group of $n\times n$ integer matrices of determinant $\pm 1$. Two decomposable forms $F,G$ as above are called $GL_{n}(\mathbb{Z})$-equivalent if
[…]
@@ 3.2. Hermite equivalence of polynomials and Hermite’s finiteness theorem
$f](\mathbf{X}):=a_{0}^{n-1}\prod_{i=1}^{n}(X_{1}+\alpha_{i}X_{2}+\cdots+\alpha_{i}^{n-1}X_{n})\in\mathbb{Z}[X_{1},\ldots,X_{n}].$
@@
Since $\mathbb{Z}$-equivalence implies $GL_{2}(\mathbb{Z})$-equivalence, it suffices to prove Theorem 3.3 for $GL_{2}(\mathbb{Z})$-equivalence. We recall the proof from BEGyRS (2023).
@@
$[g](\mathbf{X})$ $=\pm a_{0}^{n-1}\prod_{i=1}^{n}\langle\mathbf{b}_{i},\mathbf{X}\rangle\text{, where }\mathbf{b}_{i}=(\beta_{i}^{n-1},\beta_{i}^{n-2}\gamma_{i},\ldots,\gamma_{i}^{n-1})^{T}.$
@@
$=\pm c^{n-1}\prod_{i=1}^{n}\langle\mathbf{a}_{i},t(A)^{T}\mathbf{X}\rangle=\pm[f](t(A)^{T}\mathbf{X}),$
@@
In BEGyRS (2023) we gave, for every $n\geq 4$ and both for the non-monic and for the monic case, infinite collections of polynomials in $\mathbb{Z}[X]$ with degree $n$ that are Hermite equivalent but not $GL_{2}(\mathbb{Z})$-equivalent. More precisely we proved the following.
@@
    There exist infinitely many Hermite equivalence classes of properly non-monic [^1] That is, not $GL_{2}(\mathbb{Z})$-equivalent to any monic polynomial primitive  [^2] An integral polynomial is called primitive if its coefficients have greatest common divisor $1$, irreducible polynomials of degree $n$ that split into more than one $GL_{2}(\mathbb{Z})$-equivalence class.
@@
We proved Theorem 3.4 simultaneously for the cases (i) and (ii). We constructed, for every integer $n\geq 4$, an infinite parametric family of pairs $(f_{t,c}^{(n)},g_{t,c}^{(n)})$ of primitive, irreducible polynomials $f_{t,c}^{(n)}$, $g_{t,c}^{(n)}$ of degree $n$, where $c$ runs through $1$ and an infinite set of primes, and $t$ runs through an infinite set of primes with $t\neq c$ with the following properties:
(3.2) for each $n$, $f_{t,c}^{(n)},\,g_{t,c}^{(n)}$ have leading coefficient $c$ and are
properly non-monic if $c>1$;
(3.3) for each $n$, $f_{t,c}^{(n)},\,g_{t,c}^{(n)}$ are Hermite equivalent;
(3.4) for each $n$, $f_{t,c}^{(n)},\,g_{t,c}^{(n)}$ are not $GL_{2}(\mathbb{Z})$-equivalent;
(3.5) the pairs $(f_{t,c}^{(n)},g_{t,c}^{(n)})$ ($n=1,2,\ldots$) lie in different Hermite
equivalence classes.
The main steps of the proof are as follows. From the construction of $f_{t,c}^{(n)}$ and $g_{t,c}^{(n)}$ it is easy to show that (3.2) and (3.3) hold. The proof of (3.4) is more complicated. It requires the use of an irreducibility theorem of Dumas (1906), Chebotarev’s density theorem, and Dirichlet’s theorem on primes in arithmetic progressions. Finally, $f_{t,c}^{(n)}$ is so chosen that if we fix $n,c$ and let $t\to\infty$ then the absolute value of the discriminant of $f_{t,c}^{(n)}$ tends to $\infty$. By making a selection, we may assume that the discriminants of the polynomials $f_{t,c}^{(n)}$ are pairwise different. Since Hermite equivalent polynomials have the same discriminant, we obtain (3.5).
@@
We note that in our paper BEGyRS (2023) it turned out that the Hermite equivalence class of a polynomial has a very natural interpretation in terms of the so-called invariant order and invariant ideal associated with the polynomial, see Theorem 5.11 in Subsection 5.6 for more details. This fact turned out to be important in the above proofs.
Theorem 3.3 and Theorem 3.4 imply that $GL_{2}(\mathbb{Z})$-equivalence, resp. $\mathbb{Z}$-equivalence are stronger than Hermite equivalence, and hence that Hermite’s Theorem 3.2 is weaker than the most significant results of this area presented in Section 4 below.
4. Reduction theory of integral polynomials of given non-zero discriminant and of arbitrary degree
As was mentioned in the Introduction, the breakthroughs in the reduction theory due to Birch and Merriman (1972), Győry (1973), and Evertse and Győry (1991a) settled the old problem of Hermite (1857), to prove that for every given $n\geq 2$ and $D\not=0$ there are up to $GL_{2}(\mathbb{Z})$-equivalence only finitely many polynomials $f\in\mathbb{Z}[X]$ of degree $n$ and discriminant $D$, and to determine these effectively. We state the results in more detail.
4.1. The theorems of Birch and Merriman (1972), Győry (1973) and Evertse and Győry (1991a)